# Future of Work with AI Agents: Auditing Automation and Augmentation Potential across the U.S. Workforce

### v2

@@ 1 Introduction
To address these gaps, we propose a principled, survey-based framework to investigate which occupational tasks workers want AI agents to automate or augment. We look at the entire workforce that could be impacted by digital AI agents by sourcing occupational tasks from the U.S. Department of Labor’s O*NET database. Compared to occupation-level studies, task-level auditing allows us to better capture the nuanced, open-ended, and contextual nature of real-world work. Our auditing framework takes a worker-centric approach by soliciting first-hand insights from domain workers actively performing the tasks. To guide domain workers in providing well-calibrated responses, we empower them to share their experiences and articulate their reasoning through an audio-enhanced survey system. Crucially, our framework expands beyond the binary view of automation. We propose the Human Agency Scale (i.e., H1-H5), which complements the SAE L0-L5 automation levels (Committee et al., 2014) by quantifying the degree of human involvement required for occupational task completion and quality. This new scale centers human agency—a crucial factor for responsible AI agent adoption (Fanni et al., 2023)—and provides a shared language to capture the spectrum between automation and augmentation.
@@ 2.2 Emphasizing the Spectrum of Automation and Augmentation
Unlike SAE driving automation levels (Committee et al., 2014) that adopt an “AI-first” perspective, HAS provides a human-centered lens for assessing both task properties and appropriate agent development approaches. Importantly, higher HAS levels are not inherently better—different levels suit different AI roles. Tasks at H1-H2 favor automation approaches, while H3-H5 tasks benefit from augmentation strategies. Understanding the ideal level of human involvement is essential both for workers seeking to adapt their skills and for developers aiming to build context-appropriate AI agents. For instance, fully autonomous agents shall be developed for H1 scenarios, while those agents for H3 scenarios must support meaningful coordination and communication with human collaborators (Shao et al., 2024b).
@@ 2.5 Instantiating the Audit Framework to Derive WORKBank
We developed an IRB-approved, self-hosted survey interface and distributed it through crowdsourcing platforms and targeted LinkedIn outreach. Between January and May 2025, recruitment through Prolific, Upwork, and LinkedIn yielded 1,676 participants who provided 7,016 task ratings. After filtering for adequate representation ($\geq 10$ participants per occupation), we obtained assessments from 1,500 individuals across 104 occupations and calculated average worker ratings $A_{w}(t)$ and $H_{w}(t)$ for each task. We evaluate the representativeness of WORKBank by comparing its sector-level distribution with U.S. workforce data from the Bureau of Labor Statistics (Appendix D.2). As shown in Figure 3, the comparisons suggest that our database captures a broad and representative cross-section of the U.S. workforce at the sector level.
For expert assessments, we recruited 52 AI experts—PhD researchers and industry practitioners with experience in AI agent R&D. Each task was independently assessed by at least two experts, with additional reviews ensuring rating consistency (standard deviation $\leq 1$). Inter-annotator agreement, measured by Krippendorff’s $\alpha$, was 0.539 for $A_{e}(t)$ and 0.511 for $H_{e}(t)$ (see robustness analysis details in Appendix B).
@@ 3.1 Worker-centered Views on Occupational Task Automation
When discussing the absence of human qualities, workers express specific concerns about losing a “human touch” in their work, diminishing creative control, and the desire to maintain agency in decision-making. This sentiment echoes our quantitative findings from the breakdown of automation desire scores across sectors (Figure 4 a). In these sector-level breakdowns, the “Arts, Designs, and Media” sector stands out, with only 17.1% of tasks receiving positive desire ratings ($>3$ on a 5-point Likert scale). Audio responses from participants in this sector reveal nuanced opposition to automating content creation, such as: “I want it to be used for seamlessly maximizing workflow and, you know, making things less repetitive and tedious and arduous with workflow. No content creation,” “I would never use AI to like replace artists. I would be more for personal [project management] use, if anything,” “AI can be a game-changer in data architect workflow, helping to improve efficiency, accuracy and even creativity. But I create my design by myself. For research, I use AI”.
@@ 3.2 Desire-Capability Landscape for AI Agents in the Workplace
Tasks from WORKBank are broadly distributed across the landscape, with no strong correlation between $A_{w}(t)$ and $A_{e}(t)$ (Spearman $\rho=0.17$, $p<1e-6$). Overall, automation desire shows a negative correlation with both job loss concern (Spearman $\rho=-0.223$, $p<5e-11$) and enjoyment (Spearman $\rho=-0.284$, $p<3e-17$). This suggests both alignments and misalignments between worker desires and technological capabilities, with a consistent pattern: workers are less inclined to have AI agents automate tasks they enjoy or feel vulnerable about potential job loss, consistent with findings from prior literature (Armstrong et al., 2024, Gödöllei and Beck, 2023).
@@
Figure 6: Distributions on the Human Agency Scale (HAS) reveal diverse patterns of AI agent integration across occupations and underscore opportunities for human–agent collaboration. a, Comparison between worker-desired HAS levels ($H_{w}(t)$) and expert-assessed feasible HAS levels ($H_{e}(t)$) shows that workers generally prefer higher levels of human agency than what experts deem technologically necessary. b, Distribution of occupations by their dominant worker-desired HAS level shows that most occupations cluster around H3 and helps identify occupations that are at the poles of the agency spectrum. Each subplot displays the task-level distributions of worker-desired and expert-assessed HAS levels for a given occupation. Jensen-Shannon divergence (JSD) quantifies the difference between these distributions (see top 10 in Table 5). Full occupation-level results are shown in Figure 10. c, Radar plot of task characteristics based on worker ratings indicates that tasks in the H5 region are particularly associated with Interpersonal Communication. d Radar plot based on expert ratings shows that tasks in H5 are marked by strong Interpersonal Communication and Domain Expertise components.
@@ 3.3 Human Agency Scale (HAS) Spectrum
Each task in WORKBank is assigned a worker-desired and expert-rated HAS level via majority vote. Among 844 tasks, 26.9% receive matching levels from workers and AI experts. Figure 6 a shows workers generally prefer higher levels of human agency than what experts deem technologically necessary, with 47.5% of tasks fall into the lower triangle of the matrix. To quantify this divergence, we compute the Jensen-Shannon Distance (JSD) between the distributions of $H_{w}(t)$ and $H_{e}(t)$ at the occupation level. Disagreements are most pronounced in the lower HAS range as Table 5 shows that five of the ten occupations with the highest JSD scores are also those that experts rate as H1 dominant.This signals potential frictions as AI adoption progresses. Ensuring socially responsible deployment of AI agents and supporting workers currently performing low-HAS tasks warrant further scrutiny.
@@ Acknowledgements
We would like to thank Chuchu Jin and Yanzhe Zhang for their voluntary help in distributing the survey interface on Linkedin. We are grateful to Hao Zhu for database setup, to Will Held, Dora Zhao, Omar Shaikh, Yifan Mai, Sunny Yu, Zachary Robertson for their valuable feedback on the manuscript, and to all members of Stanford SALT lab for their suggestions at different stages of this project. This work would not have been possible without the thoughtful participation of the 1,500 domain workers and 52 AI experts who contributed to the study. This research is supported in part by grants from ONR grant N000142412532, and NSF grant IIS-2247357.
@@ B.1 Annotation Agreement of AI Expert Assessments
AI experts with practical R&D experience provided ratings for current automation capability $A_{e}(t)$ and feasible human-agency levels $H_{e}(t)$. To ensure high-quality annotations, we applied the following controls:
@@
      Industry practitioner (e.g., machine learning engineer or research scientist) with hands-on experience in LLMs and agentic systems.
@@
Inter-annotator agreement, measured by Krippendorff’s $\alpha$, was 0.539 for $A_{e}(t)$ and 0.511 for $H_{e}(t)$.
@@ B.2 Mixed-Effects Model Regression on Worker Responses
Table 1: Fixed-effects estimates from the mixed-effects regression predicting automation desire ratings.
To assess the extent to which workers’ automation desire ratings reflect intrinsic task properties rather than individual demographics, we fitted a linear mixed-effects model of the form
@@
where $y_{ij}$ is the automation desire rating provided by worker $i$ on task $j$, $X_{k,ij}$ are the $K$ demographic and attitude covariates (age, gender, education, experience, LLM familiarity/use, income, political affiliation, LLM-usage subtypes, and AI-attitude scales), $\beta_{k}$ their fixed effects, $u_{j}$ a task-specific random intercept, and $\varepsilon_{ij}$ the residual error. The model was estimated by REML using the statsmodels MixedLM interface in Python.
The fitted model (see Table 1) yielded the following variance-component estimates:
@@
[…]
i.e., roughly 5% of total variance in automation desire is attributable to between-task differences. This “small-to-moderate” ICC confirms that task-level properties carry a real signal in workers’ automation desire ratings.
@@
Together, these results indicate that, after controlling for a broad array of individual differences, task identity still explains a meaningful fraction of variance in automation desire. In §3, we use the average ratings for analysis.
@@ Appendix D WORKBank Statistics
As detailed in Section 2.5, we retained only occupations with at least ten worker responses from January to May 2025, yielding 1,500 individual assessments across 104 occupations. In this section, we present detailed statistics on occupational coverage and participant demographics in the WORKBank database.
@@ D.1 Full List of Included Occupations
Occupation (O*NET-SOC Title) $N$ Occupation (O*NET-SOC Title) $N$

### v3

@@ 1 Introduction
To address these gaps, we propose a principled, survey-based framework to investigate which occupational tasks workers want AI agents to automate or augment. We look at the entire workforce that could be impacted by digital AI agents by sourcing occupational tasks from the U.S. Department of Labor’s O*NET database. Compared to occupation-level studies, task-level auditing allows us to better capture the nuanced, open-ended, and contextual nature of real-world work. Our auditing framework takes a worker-centric approach by soliciting first-hand insights from domain workers actively performing the tasks. To guide domain workers in providing well-calibrated responses, we empower them to share their experiences and articulate their reasoning through an audio-enhanced survey system. Crucially, our framework expands beyond the binary view of automation. We propose the Human Agency Scale (i.e., H1-H5), which complements the SAE L0-L5 automation levels (Committee and others, 2014) by quantifying the degree of human involvement required for occupational task completion and quality. This new scale centers human agency—a crucial factor for responsible AI agent adoption (Fanni et al., 2023)—and provides a shared language to capture the spectrum between automation and augmentation.
@@ 2.2 Emphasizing the Spectrum of Automation and Augmentation
Unlike SAE driving automation levels (Committee and others, 2014) that adopt an “AI-first” perspective, HAS provides a human-centered lens for assessing both task properties and appropriate agent development approaches. Importantly, higher HAS levels are not inherently better—different levels suit different AI roles. Tasks at H1-H2 favor automation approaches, while H3-H5 tasks benefit from augmentation strategies. Understanding the ideal level of human involvement is essential both for workers seeking to adapt their skills and for developers aiming to build context-appropriate AI agents. For instance, fully autonomous agents shall be developed for H1 scenarios, while those agents for H3 scenarios must support meaningful coordination and communication with human collaborators (Shao et al., 2024b).
@@ 2.5 Instantiating the Audit Framework to Derive WORKBank
We developed an IRB-approved, self-hosted survey interface and distributed it through crowdsourcing platforms and targeted LinkedIn outreach. Between January and May 2025, recruitment through Prolific, Upwork, and LinkedIn yielded 1,678 participants who provided 7,016 task ratings. After filtering for adequate representation ($\geq 10$ participants per occupation), we obtained assessments from 1,500 individuals across 104 occupations and calculated average worker ratings $A_{w}(t)$ and $H_{w}(t)$ for each task. We evaluate the representativeness of WORKBank by comparing its sector-level distribution with U.S. workforce data from the Bureau of Labor Statistics (Appendix D.2). As shown in Figure 3, the comparisons suggest that our database captures a broad and representative cross-section of the U.S. workforce at the sector level.
For expert assessments, we recruited 52 AI experts—PhD researchers and industry practitioners with experience in AI agent R&D. Each task was independently assessed by at least two experts, with additional reviews ensuring rating consistency (standard deviation $\leq 1$). Inter‐annotator agreement, measured by Krippendorff’s $\alpha$, was 0.539 for $A_{e}(t)$ and 0.511 for $H_{e}(t)$ (see robustness analysis details in Appendix B).
@@ 3.1 Worker-centered Views on Occupational Task Automation
When discussing the absence of human qualities, workers express specific concerns about losing a “human touch” in their work, diminishing creative control, and the desire to maintain agency in decision-making. This sentiment echoes our quantitative findings from the breakdown of automation desire scores across sectors (Figure 4 a). In these sector-level breakdowns, the “Arts, Designs, and Media” sector stands out, with only 17.1% of tasks receiving positive desire ratings (>3>3 on a 5-point Likert scale). Audio responses from participants in this sector reveal nuanced opposition to automating content creation, such as: “I want it to be used for seamlessly maximizing workflow and, you know, making things less repetitive and tedious and arduous with workflow. No content creation,” “I would never use AI to like replace artists. I would be more for personal [project management] use, if anything,” “AI can be a game-changer in data architect workflow, helping to improve efficiency, accuracy and even creativity. But I create my design by myself. For research, I use AI”.
@@ 3.2 Desire-Capability Landscape for AI Agents in the Workplace
Tasks from WORKBank are broadly distributed across the landscape, with no strong correlation between $A_{w}(t)$ and $A_{e}(t)$ (Spearman $\rho=0.17$, p<1e6p<1e-6). Overall, automation desire shows a negative correlation with both job loss concern (Spearman $\rho=-0.22$, p<5e11p<5e-11) and enjoyment (Spearman $\rho=-0.28$, p<3e17p<3e-17). This suggests both alignments and misalignments between worker desires and technological capabilities, with a consistent pattern: workers are less inclined to have AI agents automate tasks they enjoy or feel vulnerable about potential job loss, consistent with findings from prior literature (Armstrong et al., 2024, Gödöllei and Beck, 2023).
@@
Figure 6: Distributions on the Human Agency Scale (HAS) reveal diverse patterns of AI agent integration across occupations and underscore opportunities for human–agent collaboration. a, Comparison between worker-desired HAS levels ($H_{w}(t)$) and expert-assessed feasible HAS levels ($H_{e}(t)$) shows that workers generally prefer higher levels of human agency than what experts deem technologically necessary. b, Distribution of occupations by their dominant worker-desired HAS level shows that most occupations cluster around H3 and helps identify occupations that are at the poles of the agency spectrum. Each subplot displays the task-level distributions of worker-desired and expert-assessed HAS levels for a given occupation. Jensen-Shannon divergence (JSD) quantifies the difference between these distributions (see top 10 in Table 7). Full occupation-level results are shown in Figure 10. c, Radar plot of task characteristics based on worker ratings indicates that tasks in the H5 region are particularly associated with Interpersonal Communication. d, Radar plot based on expert ratings shows that tasks in H5 are marked by strong Interpersonal Communication and Domain Expertise components.
@@ 3.3 Human Agency Scale (HAS) Spectrum
Each task in WORKBank is assigned a worker-desired and expert-rated HAS level via majority vote. Among 844 tasks, 26.9% receive matching levels from workers and AI experts. Figure 6 a shows workers generally prefer higher levels of human agency than what experts deem technologically necessary, with 47.5% of tasks fall into the lower triangle of the matrix. To quantify this divergence, we compute the Jensen-Shannon Distance (JSD) between the distributions of $H_{w}(t)$ and $H_{e}(t)$T at the occupation level. Disagreements are most pronounced in the lower HAS range as Table 7 shows that five of the ten occupations with the highest JSD scores are also those that experts rate as H1 dominant. This signals potential frictions as AI adoption progresses. Ensuring socially responsible deployment of AI agents and supporting workers currently performing low-HAS tasks warrant further scrutiny.
@@ Acknowledgements
We would like to thank Chuchu Jin and Yanzhe Zhang for their help in distributing the survey interface on Linkedin. We are grateful to Hao Zhu for database setup, to Will Held, Dora Zhao, Omar Shaikh, Yifan Mai, Sunny Yu, Zachary Robertson for their valuable feedback on the manuscript, and to all members of Stanford SALT lab for their suggestions at different stages of this project. This work would not have been possible without the thoughtful participation of the 1,500 domain workers and 52 AI experts who contributed to the study. This research is supported in part by grants from ONR grant N000142412532, and NSF grant IIS-2247357 and support from the Stanford Digital Economy Lab.
@@ B.1 Annotation Agreement of AI Expert Assessments
AI experts with practical R&D experience provided ratings for current automation capability $A_{e}(t)$ and feasible human‐agency levels $H_{e}(t)$. To ensure high‐quality annotations, we applied the following controls:
@@
      Industry practitioner (e.g., machine learning engineer or research scientist) with hands‐on experience in LLMs and agentic systems.
@@
Inter‐annotator agreement, measured by Krippendorff’s $\alpha$, was 0.539 for $A_{e}(t)$ and 0.511 for $H_{e}(t)$.
@@ B.2 Mixed-Effects Model Regression on Worker Responses
Table 1: Fixed‐effects estimates from the mixed‐effects regression predicting automation desire ratings.
To assess the extent to which workers’ automation desire ratings reflect intrinsic task properties rather than individual demographics, we fitted a linear mixed‐effects model of the form
@@
where $y_{ij}$ is the automation desire rating provided by worker $i$ on task $j$, $X_{k,ij}$ are the $K$ demographic and attitude covariates (age, gender, education, experience, LLM familiarity/use, income, political affiliation, LLM-usage subtypes, and AI-attitude scales), $\beta_{k}$ their fixed effects, $u_{j}$ a task‐specific random intercept, and $\varepsilon_{ij}$ the residual error. The model was estimated by REML using the statsmodels MixedLM interface in Python.
The fitted model (see Table 1) yielded the following variance‐component estimates:
@@
$\mathrm{ICC}\;=\;\frac{\widehat{\sigma}_{u}^{2}}{\widehat{\sigma}_{u}^{2}+\widehat{\sigma}_{\varepsilon}^{2}}\;=\;0.050,$
i.e., roughly 5% of total variance in automation desire is attributable to between‐task differences. This “small‐to‐moderate” ICC confirms that task‐level properties carry a real signal in workers’ automation desire ratings.
@@
We further conduct an additional analysis focusing on tasks where the task-level signal is stronger. We identified “high-signal” tasks based on two criteria: (1) low within-task variance in worker ratings (below the median variance), and (2) strong separate from the overall mean desire rating (above the median absolute deviation). This filtering retained 238 tasks. Fitting on this high-signal subset, the same linear mixed-effects model yields an ICC of 0.405. Critically, the distribution of tasks across the four zones remained remarkably stable: in the high-signal subset, 41.60% of tasks fall in the Automation “Green Light” Zone (vs. 43.13% in the full dataset), 31.09% in the Automation “Red Light” Zone (vs. 31.64%), 15.13% in the Low Priority zone (vs. 12.56%), and 12.18% in the R&D Opportunity zone (vs. 12.68%). Together, these results indicate that, after controlling for a broad array of individual differences, task identity still explains a meaningful fraction of variance in automation desire. In the main paper, we use the average ratings for analysis.
@@ Appendix D WORKBank Statistics
As detailed in Section 2.5, we retained only occupations with at least ten worker responses from January to May 2025, yielding 1,500 individual assessments across 104 occupations. In this section, we present detailed statistics on occupational coverage, participant demographics, and recruitment sources in the WORKBank database.
@@ D.1 Full List of Included Occupations
Occupation (O*NET-SOC Title) NN Occupation (O*NET-SOC Title) NN
@@ D.3 Domain Worker Demographic Information
D.4 Recruitment Sources
To obtain worker assessments, we mainly recruited domain workers through Prolific, Upwork, and LinkedIn. Table 3 reports the breakdown of recruited domain workers before data filtering. Fo expert assessments, we recruited 52 AI experts through word of mouth. These AI experts include PhD students, postdocs, professors, and industry practitioners spanning across different career stages. Table 4 reports their career stage and institutional distribution.
Recruitment Source Description # Participants
Prolific
We applied the following prescreen criteria: Employment Status (Full-Time or Part-Time);
Nationality (United States); Fluent Languages (English); Method of Work (computer or
mobile device used for $\geq$ 25% of work). We also used the "Industry" prescreen filter to
ensure participants could only select occupations matching their Prolific profile industry.
1060
Upwork We required Upwork workers to be located in the United States. 184
Word of Mouth
We distributed the survey link to LinkedIn users located in the United States and
distributed flyers within the United States.
434
Table 3: Breakdown of recruitment sources for domain workers in this study. The numbers shown reflect all participants recruited before data filtering. Breakdowns by recruitment source after data filtering are not available, as we did not link participant personal information with data points in WORKBank to preserve privacy.
Career Stage Count Breakdown
PhD Student 25
1st year (3), 2nd year (2), 3rd year (7), 4th year (6), 5th year or above (7);
Stanford University (9), Cornell University (2), University of Cambridge (2),
University of Michigan (2), Carnegie Mellon University (1),
Northwestern University (1), University of Southern California (1),
University of Texas at Austin (1), Texas A&M University (1),
Purdue University (1), Georgia Institute of Technology (1),
New York University (1), University of Illinois Urbana-Champaign (1),
The Hong Kong University of Science and Technology (1).
Postdoc 6 Stanford University (5), Harvard University (1).
Professor 6
University of Texas at San Antonio (1), Emory University (1),
Beihang University (1), University at Buffalo (1),
Texas A&M University (1), Massachusetts Institute of Technology (1).
Industry Practitioner 15
< 3 years of experience (8), 3-5 years of experience (5),
12 years of experience (2);
Google (4), Amazon (2), Apple (1), Tesla (1), Tencent (1),
Salesforce (1), xAI (1), Hugging Face (1), Other startups (3)
Table 4: Career stage and institutional distribution of 52 recruited AI experts.
@@ E.5 Top 10 Occupations By Worker-Expert Discrepancies in HAS Ratings
Skill
# Supported Tasks
in WORKBank
Average
Wage Rank
Average Required
Human Agency Rank
Organizing, Planning, and Prioritizing Work: Developing specific goals and plans to
prioritize, organize, and accomplish your work.
6 11 1
Training and Teaching Others: Identifying the educational needs of others, developing
formal educational or training programs or classes, and teaching or instructing others.
4 21 2
Staffing Organizational Units: Recruiting, interviewing, selecting, hiring, and
promoting employees in an organization.
6 6 3
Updating and Using Relevant Knowledge: Keeping up-to-date technically and
applying new knowledge to your job.
19 2 4
Developing Objectives and Strategies: Establishing long-range objectives and
specifying the strategies and actions to achieve them.
17 3 5
Guiding, Directing, and Motivating Subordinates: Providing guidance and direction
to subordinates, including setting performance standards and monitoring performance.
49 4 6
Judging the Qualities of Objects, Services, or People: Assessing the value,
importance, or quality of things or people.
47 5 7
Communicating with Supervisors, Peers, or Subordinates: Providing information to
supervisors, coworkers, and subordinates by telephone, in written form, e-mail, or in
person.
35 12 8
Providing Consultation and Advice to Others: Providing guidance and expert advice
to management or other groups on technical, systems-, or process-related topic
56 9 9
Thinking Creatively: Developing, designing, or creating new applications, ideas,
relationships, systems, or products, including artistic contributions.
117 7 10
Interpreting the Meaning of Information for Others: Translating or explaining what
information means and how it can be used.
16 13 11
Making Decisions and Solving Problems: Analyzing information and evaluating
results to choose the best solution and solve problems.
52 10 12
Monitoring Processes, Materials, or Surroundings: Monitoring and reviewing
information from materials, events, or the environment, to detect or assess problems.
40 8 13
Assisting and Caring for Others: Providing personal assistance, medical attention,
emotional support, or other personal care to others such as coworkers, customers, or
patients.
7 26 14
Getting Information: Observing, receiving, and otherwise obtaining information from
all relevant sources.
58 15 15
Monitoring and Controlling Resources: Monitoring and controlling resources and
overseeing the spending of money.
23 25 16
Analyzing Data or Information: Identifying the underlying principles, reasons, or
facts of information by breaking down information or data into separate parts.
89 1 17
Selling or Influencing Others: Convincing others to buy merchandise/goods or to
otherwise change their minds or actions.
8 16 18
Documenting/Recording Information: Entering, transcribing, recording, storing, of
maintaining information in written or electronic/magnetic form.
177 14 19
Evaluating Information to Determine Compliance with Standards: Using relevant
information and individual judgment to determine whether events or processes comply
with laws, regulations, or standards.
26 17 20
Communicating with People Outside the Organization: Communicating with peo-
ple outside the organization, representing the organization to customers, the public,
government, and other external sources. The information can be exchanged in person,
in writing, or by telephone or e-mail.
8 22 21
Processing Information: Compiling, coding, categorizing, calculating, tabulating,
auditing, or verifying information or data.
41 20 22
Estimating the Quantifiable Characteristics of Products, Events, or Information:
Estimating sizes, distances, and quantities; or determining time, costs, resources, or
materials needed to perform a work activity.
14 23 23
Performing Administrative Activities: Performing day-to-day administrative tasks
such as maintaining information files and processing paperwork.
24 18 24
Performing for or Working Directly with the Public: Performing for people or dealing
directly with the public. This includes serving customers in restaurants and stores, and
receiving clients or guests.
8 24 25
Scheduling Work and Activities: Scheduling events, programs, and activities, as well
as the work of others.
10 27 26
Establishing and Maintaining Interpersonal Relationships: Developing constructive
and cooperative working relationships with others, and maintaining them over time.
1 19 27
Table 8: Full descriptions of skills covered by the WORKBank database.